OpenStudy (anonymous):

The table below shows some values of f(x) and g(x) for different values of x. x f(x) = 2x + 24 g(x) = 5(0.5)x -2 20 -1 0 5 1 2.5 2 28 Complete the chart and determine the solution of the equation f(x) = g(x)? x = -2 x = -1 x = 1 x = 20

OpenStudy (anonymous):

f(-2) = 2(-2) + 24 f(-2) = 20 g(-2) = 5*4 g(-2) = 20 Its x = -2 because THEY BOTH EQUAL 20